OpenStudy (anonymous):

evaluate s x( (8+2x-x^2) )^1/2 dx p/s: s is for integrate

OpenStudy (mimi_x3):

I assume its an integral \[\int\limits x\sqrt{8+2x-x^{2}}dx \] Well, complete the square first; then looks like a sub

OpenStudy (anonymous):

yes. I've done the completing the square but not sure whats the next step. by part or substitution?

OpenStudy (mimi_x3):

looks like a substitution would be easier

OpenStudy (anonymous):

ok. i"ll try again. tq :)

OpenStudy (anonymous):

still dont get it. help me plez :(

OpenStudy (mimi_x3):

\[\int\limits x\sqrt{-x^{2}+2x+8} dx => \int\limits x\sqrt{-(x^{2}-2x-8)} dx => \int\limits x\sqrt{-((x^{2}-1)^{2}-(1)^{2}+8))} dx\] \[=> \int\limits x\sqrt{-(x^{2}-1)-9} dx => \int\limits x\sqrt{9-(x-1)^{2}} dx\] let u = x-1 -> x = u+1 \[=> \int\limits\left(u+1\right)\sqrt{9-u^{2}}du \]

OpenStudy (mimi_x3):

then it's trig sub

OpenStudy (mimi_x3):

\[ x= sin\theta\]

OpenStudy (anonymous):

ok mimi.i'll try. thanks again.

OpenStudy (mimi_x3):

np and for this u = 3sin\theta